Mr. flores is tiling the kitchen counter for his wife. She wants a pattern with decorative tiles mixed in with plain tiles. the plain tiles cost 60 cents a piece, and the decorative tiles cost $1.30 a piece. They budgeted $300.00 for the tile, and the job will require 430 tiles in all. What is the most number of decorative tiles that Mr. Flores can buy and still stay within budget?
Answer by stanbon(75887)

Mr. flores is tiling the kitchen counter for his wife. She wants a pattern with decorative tiles mixed in with plain tiles. the plain tiles cost 60 cents a piece, and the decorative tiles cost $1.30 a piece. They budgeted $300.00 for the tile, and the job will require 430 tiles in all. What is the most number of decorative tiles that Mr. Flores can buy and still stay within budget?
-----------------------------
Let number of plain tiles "x". Let number of decorative tiles be "y"
EQUATIONS:
1st: x+y=430
2nd: 0.60x+1.30y=300
----------
Multiply thru 2nd by 100 to get rid of the decimals:
3rd: 60x+130y = 30000
Solve 1st for x: x=430-y
Substitute into 3rd to get:
60(430-y) + 130y = 30000
25800 - 60y + 130y = 30000
70y = 4200
y=60 (# of decorative tiles)
430-y=430-60 = 370 (number of plain tiles)
---------
Cheers,
Stan H.